You just received a bonus of $2,000.

a.  Calculate the future value of $2,000​, given that it will be held in the bank for 9 years and earn an annual interest rate of 8

percent?

b.  Recalculate part (a​) using a compounding period that is​ (1) semiannual and​ (2) bimonthly.

c.  Recalculate parts (a​) and (b​) using an annual interest rate of 16 percent.

d.  Recalculate part (a​) using a time horizon of 18 years at an annual interest rate of 8 percent

.e.  What conclusions can you draw when you compare the answers in parts

​(c​) and (d​) with the answers in parts (a​) and (b​)?
Answer #1

ANSWER

a:

Future value=Present value*(1+Rate)^number of periods

= 2000*(1+8%)^9

=3998.00

b:

Semi-annual

Future value=Present value*(1+Rate)^number of periods

= 2000*(1+8%/2)^(9*2)

=4051.63

Bi-Monthly

Future value=Present value*(1+Rate)^number of periods

= 2000*(1+8%/6)^(9*6)

=4089.36

c:

Recalculate parts (a) and (b) using an annual interest rate of 8 percent.

- Annual Compounding

Future value  = 2000*(1.16)^9

= 2000*3.802

= 7605.92

-semiannual

Future value = 2000*(1+(.16/2))^(9*2)

= 6851.89

-bimonthly.

Future value = 2000*(1+(.16/6))^(9*6)

= 8283.63

d:

Recalculate part (a) using a time horizon of 18 years at an annual interest rate of 8 percent.

Future value

= 2000*(1.08)^18

= 2000*3.996019

= 7992.03

--------------------------------------------------------------

As per rules I am answering the first 4 sub parts of the question

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************